# DiPEx: Dispersing Prompt Expansion for Class-Agnostic Object Detection

### v1

@@ A.3 Baselines and Implementation Details
A.4 More Ablation Study
@@
Pseudo-labels Construction For pseudo-labeling, we utilize off-the-shelf Grounding DINO with a "generic" text prompt, which demonstrates considerable zero-shot performance, as illustrated in our pilot study. Additionally, we generate pseudo boxes by concatenating all 25 beginner nouns from WordNet [13]. We then merge the predictions from these two queries and apply Soft-NMS to eliminate overlaps. In the following section 8, we also investigate the performance of DiPEx alongside other prompt-tuning methods on the quality of pseudo-labels.
Figure 8: Study of Loss Coefficient $\gamma$
Impact on Pseudo-labeled Supervision. In this section, we investigate how the quality of pseudo-labels used for self-training impacts DiPEx’s performance, given our reliance on these training samples. Specifically, we generate pseudo-labels by querying the off-the-shelf Grounding Dino model with three different approaches: 1) using a "generic" text prompt, 2). the 25 beginner nouns from WordNet [13], and a combination of both. As shown in Table 5, the "generic" text prompt alone demonstrated considerable performance. However, we observed an improvement in Average Recall (AR) when merging the predictions generated by "generic" with the 25 beginner nouns, leading us to this study. Consequently, we use these pseudo-labels to self-train our model.
@@
Grounding-DINO @ ["generic"] 44.1 17.7 51.6 80.0 28.3 11.4 33.0 56.5
Grounding-DINO @ [25 nouns] 40.5 16.0 46.6 75.0 12.1 4.4 12.0 26.7
Grounding-DINO @ merged 51.9 24.8 61.6 85.8 19.1 7.6 19.5 42.0
DiPEx @ ["generic"] 65.5 42.7 76.2 90.3 37.0 20.5 43.3 62.4
@@
Effect of Loss Coefficient $\gamma$. Since our goal is to adequately separate child-child prompts while preserving the semantic coherence between parent-child prompts, selecting an appropriate $\gamma$ is crucial to ensure this balance. As illustrated in the bar plot below, a moderate value of $\gamma$ generally works well. As expected, a larger value (e.g., $\gamma=5$) causes child-child prompts to diverge further, damaging semantic coherence and over-regularizing the model.
Figure 9: More visualizations of class-agnostic box predictions. From the columns 1 to 4, methods are MOST [42], CutLER [49], zero-shot Grounding Dino ["Generic"] [8], and our DiPEx, respectively. The last column displays human-annotated ground truth bounding boxes from the MS-COCO dataset [31].
Figure 10: More visualizations of class-agnostic box predictions. From the columns 1 to 4, methods are MOST [42], CutLER [49], zero-shot Grounding Dino ["Generic"] [8], and our DiPEx, respectively. The last column displays human-annotated ground truth bounding boxes from the MS-COCO dataset [31].

### v2

@@ 3.2 Dispersing Prompt Expansion (DiPEx)
4 Experiments
4.1 Experimental Setup
Datasets. We conduct our experiments using two detection datasets: 1). MS-COCO [32], a large-scale object detection and instance segmentation dataset, comprising approximately 115K training images and 5K validation images across 80 classes. 2). LVIS [15] includes 2.2 million high-quality instance segmentation masks covering 1,000 class labels, resulting in a long-tailed data distribution. It consists of around 100K training images and 19.8K validation images. For class-agnostic object detection (CA-OD) setting, we merge all categories from both datasets into a single class to perform class-agnostic detection. To further validate the efficacy of DiPEx in downstream out-of-distribution object detection (OOD-OD) tasks, we evaluate our method using a rectified version of the OOD-OD benchmark. Unlike previous benchmarks [12], where samples that do not contain ID instances are manually selected and ID and OOD performance are evaluated separately, we tested our approach on the MS-COCO, which includes a mixture of both ID and OOD objects. While we followed the settings outlined in OOD-OD [12], with 20 base classes in PASCAL-VOC [13] designated as ID classes and the remaining classes treated as OOD. Our choice of dataset enhances the rigor of our evaluation by combining both ID and OOD instances, providing a more realistic assessment of our method’s real-world conditions.
Evaluation Metrics. We report results for class-agnostic object detection on both the MS-COCO and LVIS validation splits. For evaluation, we adopt official metrics from the COCO 2017 challenge. Specifically, we report average precision (AP) at IoU thresholds from 0.5 to 0.95, along with average recall (AR) across the same threshold range. We also report AR by object scale: AR@S for small, AR@M for medium, and AR@L for large objects. Details on our implementation, including those of prior works used as baselines, are provided in Appendix A.2.
Method Description AR1 AR10 AR100 AR@S AR@M AR@L AP
Selective Search [47] non-parametric 0.1 1.1 7.8 0.9 7.2 20.7 0.1
UP-DETR [8] self-training 0.2 1.4 1.4 0.0 0.2 5.8 0.1
DETReg [1] self-training 0.6 3.7 12.9 0.2 12.8 35.3 1.4
FreeSOLO [49] self-training 3.7 9.7 12.6 0.5 12.3 34.1 4.2
Exemplar-FreeSOLO [18] self-training 8.2 13.0 17.9 12.6
MOST [43] self-training 3.1 6.4 6.4 0.1 1.6 24.5 3.3
CutLER [50] self-training 6.8 19.6 32.8 13.7 37.5 60.0 29.6
Grounding DINO [“generic”] [9] zero-shot 10.3 37.8 44.1 17.7 51.6 80.0 28.3
DiPEx self-training 10.5 40.8 63.2 39.2 74.3 89.8 35.9
4.2 Main Results on Class-agnostic OD and OOD-OD
Class-agnostic OD on MS-COCO. To validate our proposed method for class-agnostic object detection, we compared it against ten different baseline methods on the MS-COCO dataset, using various metrics as reported in Table 2. We observed that non-parametric methods generally underperform compared to self-training methods due to their inability to learn and extract semantic and geometric information about objects from the dataset. In contrast, Grounding DINO, leveraging pre-trained knowledge, demonstrates strong zero-shot capabilities and achieves AR100 of 44.1% with a single text prompt, “generic”. Furthermore, CoOp, which fine-tunes prompts for Grounding DINO, enhances class-agnostic detection performance by 39.0% in AR100 compared to direct zero-shot inference. Our method, which expands the learnable prompts to a wider angular distance, surpasses all baselines by achieving the highest performance across all metrics and outperforming the leading baseline, CoOp, by 3.1% in AR100. Notably, for small objects which are challenging to localize, our method improves AR@S by 7.7% compared to CoOp, indicating that expanded prompts better capture a range of object sizes. Additionally, the proposed DiPEx achieved the highest AP of 35.9%, demonstrating the superior quality of class-agnostic detection.
Class-agnostic OD on LVIS. To further validate the efficacy of DiPEx, we conducted extensive experiments on the challenging LVIS dataset, which includes thousands of classes with a long-tail distribution. As shown in Table 3, prompt tuning methods such as CoOp [24] and CoCoOp [23] outperform zero-shot Grounding DINO when using hand-crafted prompts (e.g., “items”, “generic”, “objects”). Additionally, CoCoOp surpasses multi-object discovery baselines like CutLER [50] and HASSOD [2], by 86.7% and 51.3% in AR200, respectively. Notably, SAM [26], which was pre-trained on a vast of dataset containing millions of images and billions of masks, demonstrates strong zero-shot capabilities, surpassing all other baselines. In contrast, our proposed DiPEx outperforms SAM by 13.3% in AR200 and 21.3% in AP after only four epochs of self-training, Furthermore, DiPEx exceeds CoOp by 20.1% in AR200.
Method AR1 AR10 AR200 AR@S AR@M AR@L AP AP@S AP@M AP@L
Selective Search [47] 0.1 1.1 13.0 6.1 19.9 37.6 0.2 0.4 0.2 0.2
G-DINO [“object”] [9] 4.1 17.9 27.2 13.0 44.1 71.1 5.4 5.6 10.0 9.4
G-DINO [“generic”] [9] 3.8 16.5 20.2 6.5 34.5 67.7 9.0 4.1 17.4 30.7
G-DINO [“items”] [9] 4.0 17.8 28.0 13.9 45.3 70.7 11.6 6.3 19.6 32.0
SAM [26] 42.7 27.7 66.3 75.5 6.1
CutLER [50] 2.4 9.3 21.8 10.8 35.1 55.5 4.5 2.7 9.1 15.1
HASSOD [2] 0.2 10.6 26.9 15.6 42.2 56.9 4.9 2.8 7.9 12.2
G-DINO + CoOp [24] 4.2 19.1 40.3 23.6 63.5 83.5 14.0 8.3 23.7 32.3
G-DINO + CoCoOp [23] 4.2 19.2 40.7 24.1 63.8 84.1 13.6 8.1 22.4 30.1
DiPEx 4.3 20.1 48.4 31.9 72.6 88.2 15.2 9.3 25.3 32.8
Method Known Unknown
AP AP50 AR100 AR@S AR@M AR@L AP AP@S AP@M AP@L
Selective Search [47] 8.3 1.0 8.5 23.2 0.1 0.0 0.0 0.5
MOST [43] 5.3 0.1 1.3 22.5 0.4 0.1 0.4 1.2
CutLER [50] 34.5 15.8 41.5 62.7 5.7 2.3 6.9 13.7
VOS [12] 36.6 56.7 10.0 2.2 6.1 27.1 2.8 0.8 2.2 7.2
PROB [62] 28.2 43.8 13.2 1.9 11.2 40.3 0.9 0.6 0.9 2.1
UnSniffer [30] 35.8 55.8 20.6 11.8 19.9 34.8 2.9 1.5 3.1 5.3
G-DINO [“generic”] 46.3 59.7 43.3 18.0 52.1 82.6 12.5 6.9 17.8 25.7
DiPEx 46.3 59.7 59.9 35.8 72.9 89.7 15.7 9.7 21.8 25.2
Downstream OOD-OD on MS-COCO. To evaluate the generalization of our proposed DiPEx in out-of-distribution object detection (OOD-OD), we compared its performance on both known and unknown classes against various baselines. As shown in Table 4, the zero-shot Grounding DINO uses known class names as prompts, supplemented with a simple “generic” prompt for unknowns, outperforms all other non-VLM methods (e.g., 25.5% higher AR100 compared to CutLER [50]). This improvement stems from VLMs leveraging rich semantic knowledge from language models to better comprehend object information in images. DiPEx enhances this further by expanding text prompts in embedding space, enabling it to capture and differentiate objects of varying sizes and diverse semantics from learned classes. This approach delivers a significant performance gain, achieving a 38.3% increase in AR100 and a 25.6% in AP increase over zero-shot predictions. Furthermore, the expanded prompts can be directly applied alongside various known class vocabularies to detect unknown objects, eliminating the need for retraining.
4.3 Ablation Study and Model Analysis
We investigate the impact of various factors on prompting performance including the learnable prompt lengths, the number of expansion rounds $L$, and angular coverage achieved across rounds. To facilitate model analysis, we present the distribution of prompt logit activation and visualization of detection results. Further ablation studies refers to Section A.3.
Impact on Number of Prompts. In Figure 4, we compare the impact of prompt length $N$ for DiPEx against CoOp [24] & CoCoOp [23]. Overall, DiPEx shows consistent improvement in performance with a greater number of prompts – not merely due to quantity, but rather because a larger set fosters greater diversification, enabling the model to capture more comprehensive semantics. In contrast, CoOp’s [24] performance remains constant, while CoCoOp’s [23] performance declines, suggesting that more prompts do not necessarily guarantee enhanced performance.
Qualitative Study. In this section, we present visualized class-agnostic box predictions on images sampled from the MS-COCO dataset [32], as shown in Figure 7. The proposed DiPEx method demonstrates a superior ability to detect more bounding boxes than all baseline methods, particularly for small objects. For example, people in the distance (rows 1 and 3) and some bonsai (row 2) are missed by all baselines but successfully detected by DiPEx, showcasing its strong capability in localizing challenging small objects. For large objects, such as a motorcycle (row 3) and two people shaking hands in the near distance (row 1), DiPEx localizes them with significantly higher confidence compared to the zero-shot predictions of Grounding DINO using the prompt “generic”. Additionally, DiPEx successfully identifies objects that are not annotated in the MS-COCO ground truth, such as plates (row 1), a pillowcase (row 2), and a frame on the wall (row 2). This highlights DiPEx’s ability to identify a comprehensive set of class-agnostic objects, even those missed in human annotations.
5 Conclusion and Limitations
Acknowledgments and Disclosure of Funding
This research is partially supported by the Australian Research Council (DE240100105, DP240101814, DP230101196)
Appendix A Appendix / Supplemental Material
This supplementary material includes a comprehensive overview of related work on class-agnostic object detection, vision-language models (VLMs), and prompt tuning. Additionally, we provide detailed descriptions of baselines, implementation details for both the baselines and the proposed method, and an extensive ablation study are provided. The ablation study analyzes the impact of pseudo-labeled supervision and the effect of the hyperparameter $\gamma$. Lastly, we present more comprehensive visualizations of class-agnostic box predictions.
  • Section A.1: Related Work
  • Section A.2: Baselines and Implementation Details
  • Section A.3: More Ablation Studies
  • Figures 9 and 10: Additional Visualizations of Class-Agnostic Box Predictions
A.1 Related Work
Class-Agnostic Object Detection. Traditional bottom-up approaches [47, 27, 61, 40, 41, 6] for region proposal generation, often grapple with the constraints precision, despite high recall rates, reducing their scalability for general use in diverse environments. Recent breakthroughs in ViTs [4, 10, 39] have enabled scaling up to massive datasets for self-supervised learning, extracting both local and global semantically meaningful features. This has led to numerous methods in unsupervised object discovery and localization. LOST [45] is an early application, using a patch similarity graph and an inverse degree map to identify seed patches and extract bounding boxes. TokenCut [51] constructs an undirected graph with image tokens as nodes, applying the normalized cut algorithm [44] for foreground-background segregation. MOVE [37] builds on LOST by employing deep spectral bipartitioning, offering a more principled and effective approach. However, both LOST [45] and TokenCut [51] are limited to detecting a single object per image. MOST [43] addresses this limitation by using entropy-based box analysis (EBA) to segregate foreground tokens. Nevertheless, their performance remains sub-optimal, constrained by their limited capacity to generalize across diverse object categories. Closest to our work is MAVL [36], where they develop an MViT with late fusion strategy and use generic text queries like “all objects” to locate objects. However, their framework requires full supervision and multiple inferences with different textual prompts, yet still falls short of achieving optimal performance. In contrast, our approach eliminates the need for labels and achieves SOTA performance with one-pass inference with the non-overlapping prompts.
A.2 Baselines and Implementation Details
Implementation Details. Our code is developed on the Open Grounding-DINO framework [63], and operates on a single NVIDIA RTX A6000 GPU with 48 GB of memory. For our experiments, we choose a batch size of 8 for training, and set hyperparameter $\gamma=0.1$, $\tau_{p}=0.1$, $\tau_{c}=0.1$, $\theta=\pm 15^{\circ}$, $K=9$, $L=3$, and while adopting all remaining hyperparameters from the Open Grounding-DINO codebase. We empirically set the $\mathcal{T}_{\alpha}=75^{\circ}$ as our threshold for expansion termination. The original implementation of CoOP was developed for image classification tasks based on CLIP and supervised contrastive learning. We extend CoOP to class-agnostic object detection using pseudo labeling-based self-training, which remains consistent with our approach. All the implementation code and configurations files are provided in supplementary materials and will be publicly released upon acceptance of this work.
A.3 More Ablation Studies
@@
Pseudo-labels Construction For pseudo-labeling, we utilize off-the-shelf Grounding DINO with a “generic” text prompt, which demonstrates considerable zero-shot performance, as illustrated in our pilot study. Additionally, we generate pseudo-boxes by concatenating all 25 beginner nouns from WordNet [14]. We then merge the predictions from these two queries and apply Soft-NMS to eliminate overlaps. In the following Section 5, we also investigate the performance of DiPEx alongside other prompt-tuning methods on the quality of pseudo-labels.
@@
Grounding DINO @ [“generic”] 44.1 17.7 51.6 80.0 28.3 11.4 33.0 56.5
Grounding DINO @ [25 nouns] 40.5 16.0 46.6 75.0 12.1 4.4 12.0 26.7
Grounding DINO @ merged 51.9 24.8 61.6 85.8 19.1 7.6 19.5 42.0
DiPEx @ [“generic”] 65.5 42.7 76.2 90.3 37.0 20.5 43.3 62.4
@@
Impact on Pseudo-labeled Supervision. In this section, we investigate how the quality of pseudo-labels used for self-training impacts DiPEx’s performance, given our reliance on these training samples. Specifically, we generate pseudo-labels by querying the off-the-shelf Grounding Dino model with three different approaches: 1) using a “generic” text prompt, 2). the 25 beginner nouns from WordNet [14], and a combination of both. As shown in Table 5, the “generic” text prompt alone demonstrated considerable performance. However, we observed an improvement in Average Recall (AR) when merging the predictions generated by “generic” with the 25 beginner nouns, leading us to this study. Consequently, we use these pseudo-labels to self-train our model.
Figure 8: Study of Loss Coefficient $\gamma$
Effect of Loss Coefficient $\gamma$. To effectively separate child prompts while maintaining semantic coherence between parent and child prompts, selecting an appropriate $\gamma$ is essential. As illustrated in the bar plot below, a moderate $\gamma$ value typically yields optimal results. In contrast, a larger value (e.g., $\gamma=5$) causes child prompts to diverge more significantly, distorting semantic integrity and potentially leading to over-regularization of the model.
Figure 9: Additional visualizations of class-agnostic box predictions. Columns 1 – 4 correspond to the following methods: MOST [43], CutLER [50], zero-shot Grounding DINO [“generic”] [9], and our proposed DiPEx, respectively. The final column presents human-annotated ground truth bounding boxes from the MS-COCO dataset [32].
Figure 10: Additional visualizations of class-agnostic box predictions. Columns 1 – 4 correspond to the following methods: MOST [43], CutLER [50], zero-shot Grounding DINO [“generic”] [9], and our proposed DiPEx, respectively. The final column presents human-annotated ground truth bounding boxes from the MS-COCO dataset [32].